After a day long effort by several people the interferometer locked in LN2.
18:21 (4min) reference time, 56MHz mod depth 11dB, 6MHz mod depth 12dB
18:30 (2min) 8dB 56MHz mod depth
18:38 (2min) 5dB 56MHz mod depth
DIFFp_TX servo doesn't seem to work, I have started to increase the
gain by hand to bring UGF up from 1.3Hz to 2.0Hz (the target normaly is ~2.8Hz)
18:52 (1min) 2dB 56MHz mod depth
18:53:30 unlock probably due to an oscillation at ~14Hz. Look like
oscillation was mostly in MICH. There is also a line at ~54Hz that was
increasing.
Figure 1 unlock with low 56MHz (2dBm mod depth), had increasing lines at 14Hz and 54Hz
Figure 2 the B1s spectrum had decreasde by factor ~3 when there is 9dB less 56MHz modulation
Figure 3 this doesn't reveal a coherence between B1s and Hrec. And the RIN on B1s at 100Hz is at ~1.5e-8 1/rtHz, so that puts an upper limit on the HOM RIN.
Figure 4 Next unlock in CARM NULL 1F had a slow oscillation in SDB1 TY and BS TX ( with a period of ~30s)